OpenStudy (anonymous):

what is an equation of the line that passes through the point (4 −6) and has a slope of −3

OpenStudy (precal):

use y=mx+b your slope is -3 m=-3 y=-3x+b now we need to figure out the "b" value that is your y intercept

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (precal):

(4,-6) was given to you this means x=4 and y=-6

OpenStudy (precal):

y=mx+b -6=-3(4)+b finish solving for b

OpenStudy (anonymous):

Please solve in writing formation

OpenStudy (precal):

writing formation what is that?

OpenStudy (anonymous):

typed

OpenStudy (precal):

am I not typing???????? ok sorry don't know what to tell you the problem is 99% done, just solve for b or use the other formula provided by jdoe0001

OpenStudy (anonymous):

@Clalgee

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

@algebra1helplikern

OpenStudy (zzr0ck3r):

@abram.macedonio which part did you not understand from @precal ' s explanation?

OpenStudy (anonymous):

all dude

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (jdoe0001):

hmm looks to me pretty straightforward above

OpenStudy (anonymous):

I need some one to solve clearly

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &{\huge x_1}&{\huge y_1}\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ {\huge y_1}}}={\color{green}{ m}}(x-{\color{red}{ {\huge x_1}}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=eHPTyYbNmx4 <--- that covers it well